The length of a rectangular pool is 4m more than twice its width. The perimeter of the pool is 20m. What is the area of the pool?
Answer by hamsanash1981@gmail.com(151)

Let width be x and length = 2x+4. Perimeter = 20
2(l+b) = 20
2(x+2x+4)=20
3x +4 = 10
3x = 6
x = 2
Width = 2 and length = 8
area = 2*8 = 16 sq.m